A person pulls on a spring, stretching it 15 cm, which requires a maximum force of 750 N. Find the potential energy stored in the spring when it is stretched by 15 cm and compressed by 5 cm.


1
Expert's answer
2022-02-09T06:26:27-0500

Given:

"x_1=\rm 15\: cm=0.15\: m"

"x_2=\rm 5\: cm=0.05\: m"

"F=\rm 750\: N"


The sping constant

"k=\frac{F}{x_1}=\frac{750\:\rm N}{0.15\:\rm m}=5000\:\rm N/m"

The potential energy stored in the spring:

(a)

"E_p=\frac{kx_1^2}{2}=\rm \frac{5000*0.15^2}{2}=56.25\: J"

(b)

"E_p=\frac{kx_2^2}{2}=\rm \frac{5000*0.05^2}{2}=6.25\: J"